# Byzantine-Robust Gossip: Insights from a Dual Approach

### v1

@@ B.1.1 Notations
In the global clipping setting, we consider that, as $\boldsymbol{\tau}^{t}=\tau^{t}\textbf{1}_{\mathcal{E}}$, we can overload the clipping operators writing for $\Lambda_{h}\in\mathbb{R}^{\mathcal{E}_{h}\times d}$ and $\Lambda_{b}\in\mathbb{R}^{\mathcal{E}_{b}\times d}$,
@@
  • The unitary matrix of attacks directions on Byzantine edge
@@
B.1.2 Analysis of Global Clipping Rule (GCR) Theorem 3.7
Lemma B.1.
@@
we start from Lemma B.1
@@
Then we use Lemma B.2. ∎
Assumption B.4.
Definition 3.3 Trade-off between clipping error and Byzantine influence. By denoting
@@
Lemma B.5.
(Join control of the first order error and second order bias) Under Definition 3.3, we have that
@@
[…]
    On one side we have that,
    $\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle=\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\langle z_{i}^{t}-z_{j}^{t},z_{i}^{t}-z_{j}^{t}-\operatorname{Clip}(z_{i}^{t}-z_{j}^{t};\tau^{t})\rangle=\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}\left(\|z_{i}^{t}-z_{j}^{t}\|_{2}-\tau^{t}\right)_{+}.$
@@
    $=\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}^{2}\mathds{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}>\tau^{t}}-\tau^{t}\left(\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}\mathds{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}>\tau^{t}}-\Delta_{\infty}\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}^{2}\right).$
@@
    $\kappa^{t}\in\left[\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\mathds{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}>\tau^{t}},\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\mathds{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}\geqslant\tau^{t}}-1\right].$
@@
$\langle Z_{h}^{t}-\overline{Z^{t}_{h}},E^{t}\rangle+\eta^{-1}\langle\overline{Z^{t}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$ $\leqslant\|C_{h}^{T}Z_{h}^{t}\|_{2,\kappa^{t}}^{2}-\tau^{t}\left(\|C_{h}^{T}Z_{h}^{t}\|_{1,\kappa^{t}}-\Delta_{\infty}\|C_{h}^{T}Z_{h}^{t}\|_{1}\right)+\tau^{t}\frac{\eta|\mathcal{E}_{b}|^{2}}{n_{h}}\sum_{s=0}^{t-1}\tau^{s}$
@@
Lemma B.6.
(Control of first and 2nd order error terms together) Assume Definition 3.3 and that
@@
and that at the first order $\langle C_{g}\operatorname{Clip}(C_{g}^{T}Z_{h}^{t};\tau^{t}),\tau^{t}C_{b}U_{b}^{t}\rangle\leqslant 0$ if Byzantine aims at preventing the convergence. Thus, we will include this second-order term in the proof of the Lemma B.2.
@@
$\|E^{t}-L_{h}Z_{h}^{t}\|=\|C_{h}(\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})-\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t})+\overline{E^{t}}\|^{2}=\|C_{h}(\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})-\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t})\|^{2}+\|\overline{E^{t}}\|^{2},$
@@
$+2\langle Z_{h}^{t}-\overline{Z_{h}^{t}},E^{t}\rangle+2\eta^{-1}\langle\overline{Z^{t}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$\leqslant\eta\mu_{\max}(L_{h})\left(\|\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\|^{2}+\|\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\|^{2}\right)-2\eta\mu_{\max}(L_{h})\langle\operatorname{Clip}(C_{h}^{T}Z_{h}^{t})-C_{h}^{T}Z_{h}^{t}+C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle$
$+2\langle C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
@@
$+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$+2(1-\eta\mu_{\max}(L_{h}))\underbrace{\langle C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle}_{\leqslant\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}}+2\eta\mu_{\max}(L_{h})\underbrace{\langle C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t}),\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle}_{{\leqslant\|C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\|_{1}\Delta_{\infty}\tau^{t}}}.$
@@
$+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$+2(1-\eta\mu_{\max}(L_{h}))\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}+2\eta\mu_{\max}(L_{h})\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}-2\eta\mu_{\max}(L_{h})\|\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\|_{1}\Delta_{\infty}\tau^{t}$
@@
$+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}.\rangle$
Then, the second term is controlled using the proof of Lemma B.5.
@@
we have that, under Definition 3.3
@@
Theorem B.7.
Hence, under Definition 3.3, for
@@
We start from Lemma B.3 and use the majoration of the error from Lemma B.6:
@@
Where we used that, as $\kappa^{t}\leqslant\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\mathds{1}_{\|z_{i}^{t}-z_{j}^{t}\|\geqslant\tau^{t}}-1$, the largest term in $\|C_{h}^{T}Z_{h}^{t}\|^{2}_{2,-\kappa^{t}}$ is $(\tau^{t})^{2}$.
@@
B.1.3 Computing significant quantities
@@
Thus, assuming that, as we are in a fully-connected graph, every honest node is connected to $n_{b}$ Byzantine node, we get the result
Proposition B.9 (Simplified GCR).
@@
$\|C_{h}^{T}Z_{h}^{t}\|_{1,\kappa^{t}}\geqslant\Delta_{\infty}\|C_{h}^{T}Z_{h}^{t}\|_{1}+\frac{\eta|\mathcal{E}_{b}|^{2}}{n_{h}}\sum_{s=0}^{t}\tau^{s}$ $\iff(1-\Delta_{\infty})\|C_{h}^{T}Z_{h}^{t}\|_{1,\kappa^{t}}\geqslant\Delta_{\infty}\|C_{h}^{T}Z_{h}^{t}\|_{1,-\kappa^{t}}+\frac{\eta|\mathcal{E}_{b}|^{2}}{n_{h}}\sum_{s=0}^{t}\tau^{s}$

### v2

@@ B.1 Notations
In the global clipping setting, we consider that, as $\bm{\tau}^{t}=\tau^{t}\textbf{1}_{\mathcal{E}}$, we can overload the clipping operators writing for $\Lambda_{h}\in\mathbb{R}^{\mathcal{E}_{h}\times d}$ and $\Lambda_{b}\in\mathbb{R}^{\mathcal{E}_{b}\times d}$,
@@
  • The unitary matrix of attack on edges linking to Byzantine nodes
@@
B.2 Analysis of Global Clipping Rule (GCR) Theorem 1
Lemma 2.
@@ B.2 Analysis of Global Clipping Rule (GCR) Theorem 1
we start from Lemma 2
@@
Then we use Lemma 3. ∎
Assumption 1.
Definition 3 Trade-off between clipping error and Byzantine influence. By denoting
@@
Lemma 5.
(Join control of the first order error and second order bias) Under Definition 3, we have that
@@
    $=\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+\langle C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle,$
    Using that $C_{h}^{T}\overline{Z_{h}^{t}}=0$. On one side we have that,
    $\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle$ $=\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\langle z_{i}^{t}-z_{j}^{t},z_{i}^{t}-z_{j}^{t}-\operatorname{Clip}(z_{i}^{t}-z_{j}^{t};\tau^{t})\rangle$
    $=\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}\left(\|z_{i}^{t}-z_{j}^{t}\|_{2}-\tau^{t}\right)_{+}.$
@@
    $=\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}^{2}\operatorname{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}>\tau^{t}}$
    $\qquad-\tau^{t}\left(\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}\operatorname{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}>\tau^{t}}-\Delta_{\infty}\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\|z_{i}^{t}-z_{j}^{t}\|_{2}^{2}\right).$
@@
    $\kappa^{t}\in\left[\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\operatorname{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}>\tau^{t}},\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\operatorname{1}_{\|z_{i}^{t}-z_{j}^{t}\|_{2}\geqslant\tau^{t}}-1\right].$
@@
$\langle Z_{h}^{t}-\overline{Z^{t}_{h}},E^{t}\rangle+\eta^{-1}\langle\overline{Z^{t}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$\leqslant\|C_{h}^{T}Z_{h}^{t}\|_{2,\kappa^{t}}^{2}-\tau^{t}\left(\|C_{h}^{T}Z_{h}^{t}\|_{1,\kappa^{t}}-\Delta_{\infty}\|C_{h}^{T}Z_{h}^{t}\|_{1}\right)+\tau^{t}\frac{\eta|\mathcal{E}_{b}|^{2}}{n_{h}}\sum_{s=0}^{t-1}\tau^{s}$
@@
Lemma 6.
(Control of first and 2nd order error terms together) Assume Definition 3 and that
@@
and that at the first order $\langle C_{g}\operatorname{Clip}(C_{g}^{T}Z_{h}^{t};\tau^{t}),\tau^{t}C_{b}U_{b}^{t}\rangle\leqslant 0$ if Byzantine aims at preventing the convergence. Thus, we will include this second-order term in the proof of the Lemma 3.
@@
$\|E^{t}-L_{h}Z_{h}^{t}\|$ $=\|C_{h}(\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})-\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t})+\overline{E^{t}}\|^{2}$
$=\|C_{h}(\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})-\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t})\|^{2}+\|\overline{E^{t}}\|^{2},$
@@
$\quad+2\langle Z_{h}^{t}-\overline{Z_{h}^{t}},E^{t}\rangle+2\eta^{-1}\langle\overline{Z^{t}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$\leqslant\eta\mu_{\max}(L_{h})\left(\|\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\|^{2}+\|\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\|^{2}\right)$
$\quad-2\eta\mu_{\max}(L_{h})\langle\operatorname{Clip}(C_{h}^{T}Z_{h}^{t})-C_{h}^{T}Z_{h}^{t}+C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle$
$\quad+2\langle C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
@@
$\quad+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$\quad+2(1-\eta\mu_{\max}(L_{h}))\underbrace{\langle C_{h}^{T}Z_{h}^{t},\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle}_{\leqslant\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}}$
$\quad+2\eta\mu_{\max}(L_{h})\underbrace{\langle C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t}),\tau^{t}C_{h}^{\dagger}C_{b}U_{b}^{t}\rangle}_{{\leqslant\|C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\|_{1}\Delta_{\infty}\tau^{t}}}.$
@@
$\quad+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle\quad+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle$
$\quad+2(1-\eta\mu_{\max}(L_{h}))\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}+2\eta\mu_{\max}(L_{h})\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}$
$\quad-2\eta\mu_{\max}(L_{h})\|\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\|_{1}\Delta_{\infty}\tau^{t}$
@@
$\quad+2\langle C_{h}^{T}Z_{h}^{t},C_{h}^{T}Z_{h}^{t}-\operatorname{Clip}(C_{h}^{T}Z_{h}^{t};\tau^{t})\rangle+2\|C_{h}^{T}Z_{h}^{t}\|_{1}\Delta_{\infty}\tau^{t}+2\eta^{-1}\langle\overline{Z^{t+1}_{h}},\overline{Z^{t+1}_{h}}-\overline{Z^{t}_{h}}\rangle.$
Then, the second term is controlled using the proof of Lemma 5.
@@
we have that, under Definition 3
@@
Theorem 2.
Hence, under Definition 3, for
@@
We start from Lemma 4 and use the majoration of the error from Lemma 6:
@@
Where we used that, as $\kappa^{t}\leqslant\sum_{{(i\sim j)}\in\mathcal{E}_{h}}\operatorname{1}_{\|z_{i}^{t}-z_{j}^{t}\|\geqslant\tau^{t}}-1$, the largest term in $\|C_{h}^{T}Z_{h}^{t}\|^{2}_{2,-\kappa^{t}}$ is $(\tau^{t})^{2}$.
@@
B.2.1 On the computation of $\Delta_{\infty}$
@@
Thus, assuming that, as we are in a fully-connected graph, every honest node is connected to $n_{b}$ Byzantine node, we get the result. ∎
B.3 A simplified global clipping rule
The following proposition provides a sufficient condition for being a GCR that only requires $\kappa_{t}$ to be large enough, without any dependence on $X^{t}$.
Proposition 5 (Simplified GCR).
@@ B.3 A simplified global clipping rule
This result reads as a lower bound on the fraction $\kappa^{t}/|\mathcal{E}_{h}|$ of edges that are clipped.
@@
$\|C_{h}^{T}Z_{h}^{t}\|_{1,\kappa^{t}}\geqslant\Delta_{\infty}\|C_{h}^{T}Z_{h}^{t}\|_{1}+\frac{\eta|\mathcal{E}_{b}|^{2}}{n_{h}}\sum_{s=0}^{t}\tau^{s}$
$\iff(1-\Delta_{\infty})\|C_{h}^{T}Z_{h}^{t}\|_{1,\kappa^{t}}\geqslant\Delta_{\infty}\|C_{h}^{T}Z_{h}^{t}\|_{1,-\kappa^{t}}+\frac{\eta|\mathcal{E}_{b}|^{2}}{n_{h}}\sum_{s=0}^{t}\tau^{s}$